# FLEX: A Large-Scale Multi-Modal Multi-Action Dataset for Fitness Action Quality Assessment

### v1

@@ Abstract
With the increasing awareness of health and the growing desire for aesthetic physique, fitness has become a prevailing trend. However, the potential risks associated with fitness training, especially with weight-loaded fitness actions, cannot be overlooked. Action Quality Assessment (AQA), a technology that quantifies the quality of human action and provides feedback, holds the potential to assist fitness enthusiasts of varying skill levels in achieving better training outcomes. Nevertheless, current AQA methodologies and datasets are limited to single-view competitive sports scenarios and RGB modality and lack professional assessment and guidance of fitness actions. To address this gap, we propose the FLEX dataset, the first multi-modal, multi-action, large-scale dataset that incorporates surface electromyography (sEMG) signals into AQA. FLEX utilizes high-precision MoCap to collect 20 different weight-loaded actions performed by 38 subjects across 3 different skill levels for 10 repetitions each, containing 5 different views of the RGB video, 3D pose, sEMG, and physiological information. Additionally, FLEX incorporates knowledge graphs into AQA, constructing annotation rules in the form of penalty functions that map weight-loaded actions, action keysteps, error types, and feedback. We conducted various baseline methodologies on FLEX, demonstrating that multimodal data, multiview data, and fine-grained annotations significantly enhance model performance. FLEX not only advances AQA methodologies and datasets towards multi-modal and multi-action scenarios but also fosters the integration of artificial intelligence within the fitness domain. Dataset and code are available at here.
@@ 1 Introduction
Physical exercise—particularly weight training—serves multiple purposes, ranging from supporting healthy cognitive function to preserving musculature and bone density. However, weight training also carries a risk of physical injury, often resulting from improper form, posture, or movement. Therefore, proper coaching and maintaining correct workout form are crucial to prevent injuries and fully reap the benefits of resistance training.
@@
Figure 1: An overview of the FLEX dataset. FLEX dataset consists of a core group of 38 subjects, each performing 20 different fitness actions, repeating each action 10 times. Each action repeat was recorded from 5 viewpoints, & sEMG signals and physiological parameters (heart rate, breath rate) were simultaneously collected along with videos. The data annotations contain rich text information such as action knots (AK), error types (ET), & action feedback. (Zoom in for the best view.)
@@
Prior research has demonstrated that computer vision and AI-based solutions can assist in this area. Specifically, videos of individuals performing exercises can be analyzed by models to detect errors in workout form. In the video understanding literature, this task is commonly referred to as Action Quality Assessment (AQA)—the process of quantifying how well an action, or in this context, an exercise, is performed. To train AQA models, datasets typically consist of videos of people performing actions along with corresponding annotations such as action quality scores or error labels.
@@
However, existing fitness datasets are limited in several key aspects: 1) they cover a narrow range of exercise actions; 2) often exclude weight-loaded exercises; and 3) lack diversity in subject demographics and skill levels. To address these gaps, we introduce the FLEX dataset—a high-quality resource for exercise assessment and coaching, designed with the following desirable characteristics:
  • Multiview: High-precision markerless MoCap system used to capture video data from 5 viewpoints
  • Multimodal: A Custom wearable vest with a surface EMG system was used to record muscle activation and physiological signals synchronized with video and 3D joint data. A 3D scanner and body composition analyzer captured each subject’s point cloud and body metrics.
  • Largescale: 40+ hours of high-quality multimodal dataset—20 weight-loaded fitness actions, each performed 10 times by 38 subjects. With 7,512 samples, it is the largest dataset in AQA.
  • Weight loading: FLEX includes 20 common weight training exercises to help prevent injuries.
  • Fitness Knowledge Graph (KG): We defined key action steps and built a fitness-action knowledge graph linking action keysteps, error types, and feedback.
  • Multirepetition: Each action is performed 10 times, enabling applications like AQA-based early failure prediction in fitness and sport.
  • Subject skills-level diversity: To enable assessment of a diverse range of skill levels, we recruit 38 human subjects from 3 skill levels, namely: Novice, Amateur, and Expert.
  • Expert annotations and Quality Checks: Guided by our Fitness Knowledge Graph, professionals meticulously annotated the dataset with action keysteps, error types, AQA scores, and corrective feedback. Multilevel verification and quality checks ensured data accuracy.
FLEX dataset, with these comprehensive attributes, represents a high-quality resource that supports the development of the next generation of sophisticated AI models for fine-grained action assessment and coaching and unlocks new capabilities like Video2EMG estimation for precise muscle-level intervention and AQA-based early failure prediction in fitness and sports.
@@ 2 Related Work
Action Quality Assessment Datasets. Datasets are essential to machine learning and AQA, supporting model training, evaluation, and deployment. Since the foundational work in 2014, many high-quality AQA datasets have emerged across diverse domains, including sports pirsiavash2014assessing ; parmar2017learning , healthcaregao2014jhu ; vakanski2018data ; capecci2019kimore , fitnessparmar2022domain ; li2024egoexo ; ogata2019temporal , industrial trainingsener2022assembly101 , and generative AI evaluationchen2024gaia . Early datasets, such as MIT-Dive, MIT-Skate, UNLV-Dive, and UNLV-Vault contained only single-action categories. This was followed by multi-action and multimodal datasets such as AQA-7 and MTL-AQA (contains action scores, finegrained action class labels, verbal description of action quality).
Datasets like FineDiving split the actions into clips, representing a trend toward more detailed annotation. Furthermore, the emergence of datasets like Egoexo-Fitness and LucidAction represents a shift from crawling online videos toward in-person collection of AQA data, despite being resource-intensive. This approach guarantees control over both the dataset scale and quality. Advances in AQA datasets have contributed to developing AQA methodologies, with multimodal, fine-grained, and multi-action datasets allowing models to be more comprehensive and fine-grained, improving model generalization capabilities. However, existing datasets predominantly focus on competitive sports actions, with only a few dedicated to the fitness domain (for details see Table 1). Moreover, current fitness datasets primarily include self-loaded actions ogata2019temporal ; li2024egoexo , exhibit narrow action categories parmar2022domain , or are constrained to RGB modalities ogata2019temporal ; parmar2022domain ; li2024egoexo , thereby limiting the application and innovation.
@@
Table 1: Comparison of FLEX dataset with representative AQA datasets in terms of sample size, action types, views, modalities, annotations, & sources. V: Video, Sk: Skeleton, sE: sEMG, P: Physiological Info., S: Score, G: Grade, A: Action, F: Formation, D: Description, Fb: Feedback.
@@
Dataset Year Sample Type View Modality Annotation Domain Source
MIT-Divepirsiavash2014assessing 2014 159 1 1 V S Sport Web
MIT-Skatepirsiavash2014assessing 2014 150 1 1 V S Sport Web
UNLV-Diveparmar2017learning 2017 370 1 1 V S Sport Web
UNLV-Vaultparmar2017learning 2017 176 1 1 V S Sport Web
AQA-7parmar2019action 2019 1189 7 1 V S, A Sport Web
MTL-AQAparmar2019and 2019 1412 2 1 V S, A, D Sport Web
Waseda-Squatogata2019temporal 2019 2001 1 1 V, Sk E Fitness Camera
Fis-Vxu2019learning 2019 500 1 1 V S Sport Web
TASD-2gao2020asymmetric 2020 606 2 1 V S, A Sport Web
Rhy.Gym.zeng2020hybrid 2020 1000 4 1 V S, A Sport Web
FR-FSwang2021tsa 2021 417 1 1 V G, A Sport Web
Fitness-AQAparmar2022domain 2022 13049 3 1 V G, A Fitness Web
FineDivingxu2022finediving 2022 3000 52 1 V S, A Sport Web
LOGOzhang2023logo 2023 200 12 1 V S, A, F Sport Web
FineFSji2023localization 2023 1167 1 1 V S, A Sport Web
PaSkgao2023automatic 2023 1018 1 1 V S Sport Web
EgoExo-Fitnessli2024egoexo 2024 6131 12 6 V S, A, D Fitness Camera
LucidActiondong2024lucidaction 2024 6702 8 8 V, Sk S, A Sport MoCap
Ours FLEX 2025 7512 20 5 V, Sk, sE, P S, A, D, E, Fb Fitness MoCap
To bridge those gaps and stimulate innovation in AI-based Fitness coaching, we develop FLEX—first AQA dataset that provides multi-view videos with synchronized 3D pose, sEMG, and physiological information. Furthermore, we develop detailed, professional, comprehensive annotation rules for fitness actions that integrate multiple knowledge sources. These rules form a KG containing Action Keysteps, Error Types, and Action Advice, making the annotation rich in paired textual information.
@@
Electromyography (EMG) Datasets. EMG records electrical activity from skeletal muscles to assess their activation levels, motor coordination, and detect abnormalities. EMG devices are broadly classified into surface EMG (sEMG), which uses non-invasive skin electrodes, and needle EMG (nEMG), which involves inserting fine needles into muscle tissue alcan2023current . Owing to its non-invasive nature, existing EMG datasets predominantly utilize sEMG for applications in biomechanical analysiszhang2017extracting ; hug2019individuals ; mileti2020muscle ; jarque2020large ; santuz2022sex ; khan2020semg ; dimitrov2023hig ; hamard2023comparison ; wang2023wearable , hand gesture estimationatzori2014electromyography ; khan2020semg ; jarque2020large ; liu2021neuropose ; simpetru2022sensing ; ozdemir2022dataset ; salter2024emg2pose , etc.
Existing EMG datasets mainly link signals to motion categories, overlooking their connection to action quality. Since EMG reflects neuromuscular control and muscle function—key in fitness—activation patterns serve as objective quality markers. To address this, FLEX is the first AQA dataset to integrate EMG, using a 16-channel sEMG system fastmove at 2,000 Hz to record target muscle activity during fitness movements. sEMG and MoCap were synchronized for precise temporal alignment.
@@ 3 FLEX Dataset
We present FLEX, a multimodal, multiview dataset that addresses key limitations in existing fitness AQA datasets: limited exercise variety, lack of weight-loaded actions, and low skill diversity. To our knowledge, FLEX is the first weight training dataset with multimodal data, multiview capture, detailed exercise procedures, standardized assessment rules, and corrective feedback. This section covers data collection, annotation, quality checks, procedures, and evaluation standards.
@@ 3.1 Preliminary Setup
Figure 2: Comparison between FLEX and Egoexo-fitness datasets. EWL: equipment (barbell, dumbbell, etc.)-based weight loading; IR: level of risk of injury.
Dataset Exercises EWL IR
Fitness-AQAparmar2022domain 3 high
EgoExoli2024egoexo 12 low
Ours FLEX 20 high
@@
Bodyweight exercises require no equipment, making them simpler and less injury-prone. In contrast, weight-loaded exercises involve lifting external weights, targeting specific muscles, and following structured progression, but carry a higher injury risk if done improperly. Thus, proper technique and posture are crucial in weight training to prevent injuries. To support this, we carefully selected 20 common exercises based on the following criteria: 1) Include exercises targeting both upper and lower body muscle groups. 2) Cover complex, injury-prone joints, such as the shoulders. 3) Feature a variety of higher-risk free-weight exercises. 4) Focus on widely practiced exercises to ensure broad relevance, aligning with commonly recommended training regimens. A comparison of our dataset with Fitness-AQA and EgoExo-Fitness is provided in Figure 2.
Figure 3: Comparison of ability level diversity of the subjects in AQA datasets. N: Novice, A: Amateur, P: Professional.
Dataset N A P
Sport Domain Datasetspirsiavash2014assessing ; parmar2017learning ; xu2022finediving
Waseda-Squatogata2019temporal
Fitness-AQAparmar2022domain
EgoExo-Fitnessli2024egoexo
LucidaActiondong2024lucidaction
Ours FLEX
Diverse Subject Pool. To capture diverse skill levels, we recruited subjects across Novice, Amateur, and Expert skill brackets through surveys at our institutions and local gyms, receiving over 60 registrations from students and coaches. Each applicant completed an online questionnaire, an offline ability test, and 3D/body composition analysis to assess weight-bearing capacity and skill level. We selected 38 male subjects: 10 experts (ID: P01–P10), 8 amateurs (ID: A01–A10, excluding A06 and A09), and 20 novices (ID: N01–N10). Unlike existing datasets, ours spans all skill levels, as shown in Figure 3.
@@
Human subject safety. To ensure the safety of human subjects during data collection of risk-prone exercises, they were asked to perform exercises at 80% of their maximum weight-loaded capacity. Additionally, all subjects were covered by sports injury insurance, with on-site safety supervisors and emergency medical kits available throughout the process. Our data collection process was reviewed and approved by the Institutional Review Board (IRB) of Beijing Sport University (Approval No.: 2023313H). All participants provided written informed consent prior to participation.

### v3

# FLEX: A Largescale Multimodal, Multiview Dataset for Learning Structured Representations for Fitness Action Quality Assessment

@@ Abstract
Action Quality Assessment (AQA)—the task of quantifying how well an action is performed—has great potential for detecting errors in gym weight training, where accurate feedback is critical to prevent injuries and maximize gains. Existing AQA datasets, however, are limited to single-view competitive sports and RGB video, lacking multimodal signals and professional assessment of fitness actions. We introduce FLEX, the first large-scale, multimodal, multiview dataset for fitness AQA that incorporates surface electromyography (sEMG). FLEX contains over 7,500 multi-view recordings of 20 weight-loaded exercises performed by 38 subjects of diverse skill levels, with synchronized RGB video, 3D pose, sEMG, and physiological signals. Expert annotations are organized into a Fitness Knowledge Graph (FKG) linking actions, key steps, error types, and feedback, supporting a compositional scoring function for interpretable quality assessment. FLEX enables multimodal fusion, cross-modal prediction—including the novel Video→EMG task—and biomechanically oriented representation learning. Building on the FKG, we further introduce FLEX-VideoQA, a structured question–answering benchmark with hierarchical queries that drive cross-modal reasoning in vision–language models. Baseline experiments demonstrate that multimodal inputs, multi-view video, and fine-grained annotations significantly enhance AQA performance. FLEX thus advances AQA toward richer multimodal settings and provides a foundation for AI-powered fitness assessment and coaching. Dataset and code are available at https://github.com/HaoYin116/FLEX. Link to Project page.
@@ 1 Introduction
Figure 1: An overview of the FLEX dataset. FLEX dataset consists of a core group of 38 subjects, each performing 20 different fitness actions, repeating each action 10 times. Each action repeat was recorded from 5 viewpoints, & sEMG signals and physiological parameters (heart rate, breath rate) were simultaneously collected along with videos. The data annotations contain rich text information such as action keysteps (AK), error types (ET), & action feedback. (Zoom in for the best view.)
@@
Weight training supports cognition, muscle, and bone health but carries injury risks from poor form. Proper coaching and correct technique are essential to prevent injuries and maximize benefits. Prior research has demonstrated that computer vision and AI-based solutions can assist in this area. Specifically, videos of individuals performing exercises can be analyzed by models to detect errors in workout form. In the video understanding literature, this task is commonly referred to as Action Quality Assessment (AQA)—the process of quantifying how well an action, or in this context, an exercise, is performed. To train AQA models, datasets typically consist of videos of people performing actions along with corresponding annotations such as action quality scores or error labels.
@@
However, existing fitness datasets are limited in several key aspects: 1) they cover a narrow range of exercise actions; 2) often exclude weight-loaded exercises; 3) lack diversity in subject skill levels; and 4) lack of explicit reasoning structure. To address these gaps, we introduce the FLEX dataset—a high-quality resource for exercise assessment and coaching. Our main contributions are:
  1. 1.
    FLEX Dataset designed to have the following desirable characteristics: Multiview—five-view markerless MoCap video; Multimodal—synchronized surface EMG, RGB video, 3D joints, point cloud, and body metrics; Large-scale—7,512 samples (40+ hours, 38 subjects, 20 exercises × 10 repetitions); Weight-loaded—20 common training exercises to study injury prevention; Multi-repetition—10 repetitions per action enabling early-failure prediction; Skill-diverse—38 subjects spanning Novice, Amateur, and Expert levels. FLEX overview provided in Figure 1.
  2. 2.
    Structured Annotations and Fitness Knowledge Graph. Guided by domain experts, each sequence is annotated with action key steps, error types, and quality scores. These annotations are organized into a Fitness Knowledge Graph linking actions, steps, errors, muscles, and corrective feedback, and are paired with a compositional scoring function that aggregates penalties to yield interpretable action-quality metrics. Whereas existing AQA datasets typically provide quality scores without explicit reasoning structure, FLEX naturally encodes reasoning traces through this graph, enabling models to explain quality judgments by tracing from global scores to key-step penalties, specific errors, and corresponding corrective feedback.
  3. 3.
    FLEX-VideoQA Benchmark. Leveraging the FKG, we create FLEX-VideoQA, a fine-grained video question–answering benchmark with hierarchical questions ranging from coarse action recognition to fine-grained error diagnosis and causal feedback generation. This benchmark evaluates a model’s ability to perform multi-hop, compositional reasoning over multimodal signals.
  4. 4.
    Multimodal, Cross-modal, and Biomechanically-oriented Representation Learning. FLEX supports diverse representation-learning paradigms—multimodal fusion, cross-modal prediction, and biomechanically grounded learning—by pairing synchronized RGB video, 3D pose, surface EMG, and physiological signals. Baseline models for action quality assessment and Video→EMG prediction demonstrate that visual features can be trained to infer hidden muscle activation, while experiments with FLEX-VideoQA showcase cross-modal reasoning in vision–language models.
These contributions establish FLEX as a comprehensive platform for studying fine-grained human action, enabling multimodal & biomechanically grounded representation learning & reasoning.
@@ 2 Related Work
Table 1: Comparison of FLEX dataset with representative AQA datasets in terms of sample size, action types, views, modalities, annotations, skills level coverage & sources. V: Video, Sk: Skeleton, sE: sEMG, P: Physiological Info., S: Score, G: Grade, A: Action, F: Formation, D: Description, Fb: Feedback. N: Novice, Am: Amateur, Ex: Expert.
@@
Dataset Sample Type View Modality Annotation Skill Levels Domain Source
MIT-Dive Pirsiavash et al. (2014) 159 1 1 V S Ex Sport Web
UNLV-Dive Parmar & Tran Morris (2017) 370 1 1 V S Ex Sport Web
AQA-7 Parmar & Morris (2019a) 1189 7 1 V S, A Ex Sport Web
MTL-AQA Parmar & Morris (2019b) 1412 2 1 V S, A, D Ex Sport Web
Waseda-Squat Ogata et al. (2019) 2001 1 1 V, Sk E N Fitness Camera
TASD-2 Gao et al. (2020) 606 2 1 V S, A Ex Sport Web
Rhy.Gym. Zeng et al. (2020) 1000 4 1 V S, A Ex Sport Web
FR-FS Wang et al. (2021) 417 1 1 V G, A Ex Sport Web
Fitness-AQA Parmar et al. (2022) 13049 3 1 V G, A Am Fitness Web
FineDiving Xu et al. (2022) 3000 52 1 V S, A Ex Sport Web
LOGO Zhang et al. (2023) 200 12 1 V S, A, F Ex Sport Web
FineFS Ji et al. (2023) 1167 1 1 V S, A Ex Sport Web
EgoExo-Fitness Li et al. (2024) 6131 12 6 V S, A, D N, Am Fitness Camera
LucidAction Dong et al. (2024) 6702 8 8 V, Sk S, A Am, Ex Sport MoCap
[…]
Action Quality Assessment Datasets. Datasets are essential to machine learning and AQA, supporting model design and development. Many high-quality AQA datasets have emerged across diverse domains, including sports Pirsiavash et al. (2014); Parmar & Tran Morris (2017), healthcare Gao et al. (2014); Vakanski et al. (2018); Capecci et al. (2019), fitness Parmar et al. (2022); Li et al. (2024); Ogata et al. (2019), industrial training Sener et al. (2022), and generative AI evaluation Chen et al. (2024). Early datasets, such as MIT-AQA and UNLV-AQA contained only single-action categories. This was followed by multi-action and multimodal datasets such as AQA-7 and MTL-AQA (contains action scores, fine-grained action class labels, verbal description of action quality). Datasets like FineDiving split the actions into clips, representing a trend toward more detailed annotation.
Emergence of datasets like Egoexo-Fitness and LucidAction represents a shift from crawling online videos toward in-person collection of AQA data, despite being resource-intensive. This approach guarantees control over both the dataset scale and quality. Advances in AQA datasets have contributed to developing AQA methodologies, with multimodal, fine-grained, and multi-action datasets allowing models to be more comprehensive and fine-grained, improving model generalization capabilities. However, existing datasets predominantly focus on competitive sports actions, with only a few dedicated to the fitness domain (for details see Table 1). Moreover, current fitness datasets primarily include self-loaded actions Ogata et al. (2019); Li et al. (2024), exhibit narrow action categories Parmar et al. (2022), or are constrained to RGB modalities Ogata et al. (2019); Parmar et al. (2022); Li et al. (2024), thereby limiting the application and innovation. To bridge those gaps and stimulate innovation in AI-based Fitness coaching, we develop FLEX—first AQA dataset that provides multi-view videos with synchronized 3D pose, sEMG, and physiological information. Furthermore, we develop detailed, professional, comprehensive annotation rules for fitness actions that integrate multiple knowledge sources. These rules form a KG containing Action Keysteps, Error Types, and Feedback, making the annotation rich in paired textual information.
@@
Electromyography (EMG) Datasets. EMG records electrical activity from skeletal muscles to assess their activation levels, motor coordination, and detect abnormalities. EMG devices are broadly classified into surface EMG (sEMG), which uses non-invasive skin electrodes, and needle EMG (nEMG), which involves inserting fine needles into muscle tissue ALCAN & ZİNNUROĞLU (2023). Owing to its non-invasive nature, existing EMG datasets predominantly utilize sEMG for applications in biomechanical analysis Zhang et al. (2017); Hug et al. (2019); Jarque-Bou et al. (2020); Khan et al. (2020); Dimitrov et al. (2023); Wang et al. (2023), hand gesture estimation Atzori et al. (2014); Liu et al. (2021); Ozdemir et al. (2022); Salter et al. (2024), etc. Existing EMG datasets mainly link signals to motion categories, overlooking their connection to action quality. Since EMG reflects neuromuscular control and muscle function—key in fitness—activation patterns serve as objective quality markers. FLEX is the first AQA dataset to integrate EMG, using a 16-channel sEMG system FASTMOVE (2024a) at 2 KHz to record target muscle activity during fitness movements. sEMG & MoCap were synchronized for precise temporal alignment.
@@ 3 FLEX Dataset
We present FLEX, a dataset that addresses key limitations in existing fitness AQA datasets, such as limited exercise variety, absence of weight-loaded actions, and low skill diversity. To our knowledge, FLEX is the first weight training dataset to include multiview multimodal recordings, detailed exercise procedures, standardized assessment rules, and corrective feedback. This section outlines data collection, annotation, quality checks, procedures, and evaluation standards.
FLEX dataset advances action quality assessment beyond traditional video benchmarks by combining five synchronized RGB views with 3D motion-capture, surface EMG, and physiological signals. It includes 20 weight-training actions performed by 38 subjects of varying skill levels, each repeated 10 times. By pairing video with sEMG—capturing hidden muscle activity—FLEX enables learning not only what movements are performed but also how they are executed internally.
To make this rich sensor data useful, we built a Fitness Knowledge Graph (FKG) that encodes the structure of each exercise. Every action is decomposed into ordered key steps—such as the setup, eccentric phase, and concentric phases of a squat. Within each step, annotators list common errors like knee valgus or back rounding, connect them to the primary and secondary muscle groups involved, and attach recommended feedback for correction. These nodes and edges form a hierarchy of actions, steps, errors, muscles, and advice. As a result, every clip in FLEX is not just tagged with a single action quality label or score, but grounded in a web of relationships that describe both the mechanics of the motion and the typical ways it can fail and actionable feedback to rectify it.
We formalize this structure into a compositional scoring function. Each key step carries a weight reflecting its importance to overall performance. Within a step, each error type has a penalty coefficient scaling with the observed severity of that error. The final quality score is computed by summing weighted step scores and subtracting accumulated penalties, producing a number that reflects the layered decomposition of the action. Since penalties for individual errors roll up into step scores, and step scores roll up into the global score, the metric mirrors the graph’s hierarchy and provides interpretable pathways from low-level mistakes to the final assessment.
@@ 3.1 Preliminary Setup
Figure 2: Comparison between FLEX and existing SOTA fitness datasets. EWL: equipment (barbell, dumbbell, etc.)-based weight loading; RI: level of risk of injury.
Dataset Exercises EWL RI
Fitness-AQA Parmar et al. (2022) 3 high
EgoExo Li et al. (2024) 12 low
[…]
@@
Bodyweight exercises require no equipment, making them simpler and less injury-prone. In contrast, weight-loaded exercises involve lifting external weights and carry a higher injury risk if done improperly. Thus, proper technique and posture are crucial in weight training to prevent injuries. To support this, we carefully selected 20 common exercises based on the following criteria: 1) Include exercises targeting both upper and lower body muscle groups. 2) Cover complex, injury-prone joints, such as the shoulders. 3) Feature a variety of higher-risk free-weight exercises. 4) Focus on widely practiced exercises to ensure broad relevance, aligning with commonly recommended training regimens. A comparison of our dataset with Fitness-AQA and EgoExo-Fitness is provided in Figure 2.
Diverse Subject Pool. To capture diverse skill levels, we recruited subjects across Novice, Amateur, and Expert skill brackets through surveys at our institutions and local gyms, receiving over 60 registrations from students and coaches. Each applicant completed an online questionnaire, an offline ability test, and 3D/body composition analysis to assess weight-bearing capacity and skill level. Following prior work Li et al. (2024), we selected a total of 38 subjects: 10 experts (ID: P01–P10), 8 amateurs (ID: A01–A10, excluding A06 and A09), and 20 novices (ID: N01–N10). Unlike existing datasets, ours spans all skill levels, as shown in Table 1 and Figure 2.
@@
Human subject safety. Details on ensuring human subject safety provided in the Appendix B.